Is there a different value for the answers among lim(3+x), lim3+x, and lim(x+3)? . For all three values, lim x(arrow sign)2.

Answer by stanbon(75887)

Is there a different value for the answers among
lim(3+x) as x goes to 2 = 5
lim3+x as x goes to 2 = 3 + x

lim(x+3) as x goes to 2 = 5
Cheers,
Stan H.